Consider the following data:

62 61 83 92 67 89 83 50 80 69 95 50 73 71 73 88 66 89 73 51 58 86 73 57 67 80 59 86 62 95 63 88 59 64 54 78 84 77 71 82 82 83 81 97 60 85 61 55 80 72

Make frequency distribution using 5 classes and find the following:

Arithmetic Mean

Median

Mode

All quartiles

3rd Deciles

46th Percentile

Coefficient of Skewness

Semi Inter Quartile Range (SIQR)

Standard Deviation

Mean Deviation

Variance

Sketch Histogram

Sketch OGIVE


1
Expert's answer
2021-05-26T14:08:06-0400


Arithmetic Mean:

"\mu=\frac{\sum x_i}{n}=\frac{3664}{50}=73.28"


Put numbers in increasing order:

"50,50,51,54,55,57,58,59,59,60,61,61,62,62,63,64,66,67,67,69,71,71,72,"

"73,73,73,73,77,78,80,80,80,81,82,82,83,83,83,84,85,86,86,88,88,89,89,"

"92,95,95,97"


Median (the number in the middle of the list):

"m=73"


Mode (the number which appears most often in a set of numbers):

"M=73"


Quartiles:

1st quartile is number on the (N/4)th place:

"Q_1=61"

2nd quartile is median:

"Q_2=m=73"

3rd quartile is number on the (3N/4)th place:

"Q_3=84"


3rd Decile:

"D_3=value\ of\ 3(50+1)/10=1.7th\ number", which is 0.7 of the way between 1st and 2nd numbers (50 and 50).

"D_3=50"


46th Percentile:

The 46th percentile is the lowest number in a data set that is greater than 46% of the numbers:

index="0.46\cdot50=23"

"Q_{46}=N_{24}=73"


Coefficient of Skewness:

"S_k=3\frac{\mu-m}{\sigma}"

"\sigma=\sqrt{\frac{\sum (x_i-\mu)^2}{n}}=\sqrt{\frac{7971.9216}{50}}=12.63"

"S_k=\frac{72.28-73}{12.63}=0.022"


Semi Inter Quartile Range (SIQR):

"SIQR=\frac{Q_3-Q_1}{2}=\frac{84-61}{2}=11.5"


Standard Deviation:

"\sigma=12.63"


Mean Deviation:

"MD=\frac{\sum |x_i-\mu|}{n}=\frac{555.12}{50}=11.10"


Variance:

"V=\sigma_2=12.63^2=159.52"


Histogram:



OGIVE: